A Carnot engine has an efficiency of 50% when it sinks temperature is at 27 degree celcius calculate the change in source temperature for increasing its efficiency to 60%

1
Expert's answer
2022-04-17T17:07:05-0400

"\eta=1-\frac{T_2}{T_1},"

"0.5=1-\frac{300}{T_1},"

"T_1=600~K,"

"0.6=1-\frac{300}{T_3},"

"T_3=750~K,"

"\u2206T=T_3-T_1=150~K."